OpenStudy (anonymous):

find f from f'' t = 9 divided by square root t ...... f(4) =27 and f'(4) = 10

OpenStudy (anonymous):

\[f \prime \prime = \frac{9}{\sqrt{t}} = 9 \times t ^{?}\] can you figure this out?